RESONANT MULTIPHOTON FRAGMENTATION SPECTROSCOPY OF NIOBIUM DIMER CATION IN A REFLECTRON TIME-OF-FLIGHT (TOF) MASS SPECTROMETER AND DENSITY FUNCTIONAL CAL-CULATIONS.

\underline{M. AYDIN}, Department of Chemistry, Faculty of Art and Sciences, Ondokuz Mayis University, 55139, Samsun, Turkey; JOHN R. LOMBARDI, Department of Chemistry and Center for Analysis of Structures and Interfaces (CASI), The City College of New York (CCNY), New York, NY 10031.

Resonant multiphoton fragmentation spectra of niobium dimer cation (Nb_{2}^{+}) have been obtained by utilizing laser vaporization of an Nb metal target. Ions are mass-selected with a time-of-flight mass spectrometer followed by a mass gate, then fragmented with a pulsed dye laser, and resulting fragment ions are detected with a second time-of-flight reflectron mass spectrometer and multichannel plate. Photon resonances are detected by monitoring ion current as a function of fragmentation laser wavelength. A rich, but complex spectrum of the cation is obtained. The bands display a characteristic multiple structure, which may be interpreted as involving transitions from the X^{4}Σ_{g}^{−} ground state to several excited states. The second order spin orbit splitting in the ground sate of Nb_{2}^{+} was measured to be about 142 ± 5 cm^{−1}. In addition various DFT were performed to calculate the quartet- and doublet-electronic energy levels of the Nb_{2}^{+} and the force constants (k) and inter nuclear distances (R) of the neutral and ionic dimer molecules from the first-row to third-row transitions metals at their ground states. R dependence of the logarithmic values of these calculated force constants, ln(k), provided an analogous linear equations for each set of data such as: ln(k(anion)) = 6.9223-2.8994*R, ln[k(neutral)] = 6.7887-2.7985*R, ln(k(cation)) = 5.9952-2.4163*R for the first row transitions metals; ln(k(anion)) = 7.0685-2.5864*R, ln(k(neutral)) = 7.1295-2.5990*R, ln(k(cation)) = 7.4209-2.7656*R for the second row transitions metals and ln(k(anion)) = 7.2073-2.5432*R, ln(k(neutral)) = 7.0825-2.4676*R, ln(k(cation)) = 7.4625-2.6191*R for the third row transitions metals.